# AB-UPT: Scaling Neural CFD Surrogates for High-Fidelity Automotive Aerodynamics Simulations via Anchored-Branched Universal Physics Transformers

### v3

@@ 5 Limitations and future work
AB-UPT employs expressive self-attention among anchor tokens, which exhibits quadratic complexity. We show that on high-fidelity CFD simulations a small set of anchor tokens is sufficient to obtain good results on currently available public datasets. However, using more anchor tokens can improve accuracy further, leading to (quadratically) increasing computational demands. Contrary, the computational demands of models with linear complexity would only increase linearly.
@@ 6 Conclusion
Appendix A Extended results
A.1 Training without query tokens
Training with loss on Loss when evaluated on
(i) 0.13574 0.13624 $+0.37\%$ 11.5s
(ii) 0.13634 0.13632 $\mathbf{-0.01\%}$ 19.4s
(iii) 0.13462 0.13468 $\mathbf{+0.04\%}$ 19.4s
A.2 Model design on DrivAerML
@@
A.3 Benchmarking AB-UPT against DoMINO
Table 9: Relative L2 errors (in %) of surface pressure $\bm{p}_{s}$, wall shear stress $\bm{\tau}$ per $x,y,z$, dimension, volume velocity $\bm{u}$ per $x,y,z$ dimension, and volume pressure $\bm{p}_{v}$. Lower percentage values indicate better performance in terms of L2 error. We provide the results for DrivAerML. AB-UPT outperforms other neural surrogate models, often by quite a margin. The results for DoMINO are taken from (Ranade et al., 2025)*.
Surface Volume
$\bm{p}_{s}$ $\bm{\tau}_{x}$ $\bm{\tau}_{y}$ $\bm{\tau}_{z}$ $\bm{p}_{v}$ $\bm{u}_{x}$ $\bm{u}_{y}$ $\bm{u}_{z}$
DoMINO* 15.05 21.24 30.2 33.59 21.93 23.97 50.25 45.67
AB-UPT 3.76 5.35 3.65 3.63 6.29 4.43 3.04 2.61
We trained AB-UPT with the same train/test split as provided in (Ranade et al., 2025), and report median evaluation metrics over 5 training runs. All other hyperparameters for AB-UPT remained consistent with those reported in Appendix B. The numbers for DoMINO are taken directly from (Ranade et al., 2025) and we compare to the same metrics.
A.4 Full results: benchmarking AB-UPT against other neural surrogate models
Table 10: Relative L2 errors (in %) of surface pressure $\bm{p}_{s}$, volume velocity $\bm{u}$, volume vorticity $\bm{\omega}$, wall shear stress $\bm{\tau}$, and volume pressure $\bm{p}_{v}$. Lower percentage values indicate better performance in terms of L2 error. We provide the results for ShapeNet-Car, AhmedML, and DrivAerML. AB-UPT outperforms other neural surrogate models, often by quite a margin. For ShapeNet-Car, only surface pressure and volume velocity are available.
ShapeNet-Car AhmedML DrivAerML
PointNet 12.09 3.05 8.51 5.43 67.30 10.07 6.12 23.63 28.12 880.9 41.85 31.23
GRAPH U-NET 10.33 2.49 6.42 4.14 53.53 7.21 5.17 16.13 17.98 380.4 27.85 20.52
GINO 13.28 2.53 7.90 6.23 71.81 8.18 8.10 3.03 40.58 98.5 21.71 44.90
LNO 9.05 2.29 12.91 7.56 72.05 11.44 8.47 20.51 23.27 173.0 36.44 27.02
UPT 5.60 1.40 4.25 2.73 15.03 5.80 3.10 7.44 8.74 90.2 12.93 10.05
OFormer 7.05 1.61 3.55 2.21 7.78 4.10 2.27 4.48 6.64 60.4 8.92 7.10
Transolver 6.46 1.62 3.45 2.05 8.22 4.00 2.16 4.81 6.78 38.4 8.95 7.74
Transformer 4.86 1.17 3.41 2.09 6.76 4.03 2.16 4.35 6.21 47.9 8.26 6.27
AB-UPT 4.81 1.16 3.01 1.90 6.52 3.88 1.98 3.82 5.93 35.1 7.29 6.08
A.5 Benchmarking AB-UPT against Erwin
Table 11: Benchmarking our adaption of Erwin to jointly model surface and volume data on ShapeNet-Car. Relative L2 errors (in %) of surface pressure $\bm{p}_{s}$, and volume velocity $\bm{u}$. Erwin (reimpl.) with ball size 8192 would be roughly equivalent to the Transformer row.
Ball size $\bm{p}_{s}$ $\bm{u}$
Linear complexity transformers
LNO - 9.05 2.29
OFormer - 7.05 1.60
Transolver - 6.46 1.62
Erwin (reimpl.) 512 5.46 1.37
Erwin (reimpl.) 1024 5.38 1.37
Erwin (reimpl.) 2048 5.35 1.27
Erwin (reimpl.) 4096 5.31 1.25
Quadratic complexity transformers
Transformer 8192 4.86 1.17
AB-UPT 4096 4.81 1.16
Next, we compare our reimplementation of Erwin on DrivAerML in the default setting considered in the main paper (16K/16K surface/volume points) and show results in Figure 10. We observe that ball attention imposes an overhead, most likely from memory restructuring operations to structure tokens for efficient attention computations. This overhead becomes negligible with more inputs and ball attention becomes faster than (branched) self-attention. As there is a significant accuracy gap between AB-UPT trained on 16K/16K points and Erwin (reimpl.) trained on 16K/16K points we hypothesize that more sophisticated training protocols are necessary to train ball tree attention models to sufficient quality (e.g., hierarchical architectures, different ball size to input token ratios, …). As the design space thereof is large and fair comparisons to AB-UPT become more and more difficult, we leave exploration of this direction to future work.
Finally, we compare AB-UPT against the original Erwin architecture (including three layer message passing and distance-based attention biases) in the setting used in the original paper (Zhdanov et al., 2025). Table 12 confirms that our reimplementation obtains sensible results and that AB-UPT also performs best with this protocol.
$\bm{p}_{s}$ MSE
Erwin-S 15.85
Erwin-M 15.43
Erwin (reimpl.) 15.17
AB-UPT 13.25
A.6 Anchor attention versus k-NN interpolation
Anchor attention updates the representation of a query token based on the anchor tokens, without influencing them. Additionally, we sample random points in space, so each query token most likely has a relatively close-by anchor token. To validate the effectiveness of anchor attention, we compare our approach against a baseline of a k-NN interpolation (Qi et al., 2017b), where we interpolate the anchor tokens after the full forward pass to query locations via a weighted sum of the 3 nearest neighbors of each query location where the weights are based on the distance to each neighbor.
Table 13 shows that anchor attention, although only trained on a reduced resolution (e. g., trained with 16K anchor points) can produce high-fidelity predictions also for query points (e. g.., up to 140 million points). It is significantly better than a simple k-NN interpolation as the model learned non-linear interactions between anchor points during training, which it can transfer “zero-shot”-like to query points, even though the query points do not influence the anchor points. This underlines the effectiveness of our approach and confirms that our model learns more sophisticated prediction schemes than linear interpolation.
AhmedML
$\bm{p}_{s}$ $\bm{u}$ $\bm{\omega}$
k-NN interpolation 0.00703 0.06641 0.5748
Anchor attention 0.00276 0.00657 0.0397
A.7 Extended comparison against Transolver
In Transolver (Wu et al., 2024), the quadratic attention is limited to a subset of tokens by slicing the input domain into a reduced domain (i.e., physics tokens). Throughout the experiments in the main paper, we use 512 slices. In this section, we investigate the impact of this hyperparameter in terms of performance (test loss), memory consumption, and runtime. Figure 11 shows that increasing the number of slices slightly improves test loss at the cost of additional memory and runtime requirements. As the slicing mechanism is implemented via slow and memory-heavy torch.einsum operations (following the original implementation [^4] https://github.com/thuml/Transolver/blob/main/Car-Design-ShapeNetCar/models/Transolver.py), it requires much more memory than AB-UPT, which leverages the constant memory complexity of FlashAttention (Dao et al., 2022) (which results in constant memory complexity for the attention layer, but the overall memory complexity is still linear). For larger problem sizes (e.g., DrivAerML), the memory bottleneck becomes significant and requires memory-saving techniques such as activation checkpointing to fit onto a single GPU, which in turn increases runtime. Note that we improved speed and memory consumption of the original Transolver implementation by using FlashAttention for the self-attention between slice tokens and employing mixed-precision training. Further optimizations (e.g., specialized hardware implementations of the transolver attention) are certainly possible but are beyond the scope of this work. Note that we also do not optimize AB-UPT beyond the use of FlashAttention555torch.nn.functional.scaled_dot_product_attention using PyTorch 2.4.1 with CUDA 12.4 with mixed-precision training. Also AB-UPT could be optimized further, e.g., by employing FlashAttention-3 (Shah et al., 2024).
A.8 Geometry branch design
@@
Appendix B Experimental setup
B.1 Datasets
ShapeNet-Car
@@ B.1 Datasets
DrivAerML
B.2 Evaluation metrics
[…]
B.3 AB-UPT
We choose the architecture such that surface/volume tokens traverse $12$ Transformer blocks. The first block in each branch is a cross-attention block to the output of the geometry branch. The next $8/8/4$ blocks consist of interleaved cross-attention and self-attention blocks, where cross-attention blocks exchange information between branches by using tokens of the other branch as keys and values. The self-attention blocks operate only within their branch. The last $3/3/7$ blocks are also self-attention blocks. Afterwards, a linear layer decodes the token representation into 4 surface variables (1 pressure dimension, $3$ wall shear stress dimensions) and $7$ volume variables ($1$ pressure dimension, 3 velocity dimension, $3$ vorticity dimensions).
We use a standard ViT (Dosovitskiy et al., 2021)-style pre-norm Transformer blocks where we additionally add Rotary Position Embeddings (RoPE(Su et al., 2024) to all blocks. For the message passing of the supernode pooling layer, we the positions relative to the supernode and the magnitude of the distance as input to the message passing. This is then followed by embedding the position of the supernode, concatenating it with the output of the message passing and down-projecting it to the original dimension in order to integrate absolute positions after the message passing. This is in contrast to Alkin et al. (2024a), where the concatenation of the positions of the two connected nodes are used directly for message passing.
B.4 Baseline models
GINO
Graph U-Net
LNO
The Latent Neural Operator (LNO) is a transformer-based neural operator that operates in a low-dimensional latent space, rather than the high-dimensional input/geometry space. LNO consists of an embedding layer that maps the raw input features/function to high-dimensional latent representations. A physics-cross-attention block that maps the input embedding to a low-dimensional number of latent tokens. Followed by several Transformer blocks that process these latent tokens. Finally, a decoder that maps the processed latent tokens back to the input space, predicting the PDE solution. We use the standard ViT-tiny configuration with $12$ Transformer blocks, a channel size of $192$, $3$ self-attention heads, and an up-projection ratio of $4$. Our implementation is based on (Wang & Wang, 2024), but adapted for the 3D aerodynamics. [^7] https://github.com/L-I-M-I-T/LatentNeuralOperator/blob/main/LNO-PyTorch/ForwardProblem/module/model.py
PointNet
OFormer
The OFormer (Li et al., 2023a) is a transformer-based encoder-decoder architecture which leverages linear attention mechanism (Cao, 2021) to deal with high-dimensional problems/functions. The raw input features are first embedded with RoPE positional information into tokens. Next, several Transformer blocks are applied. Finally, a decoder (i.e., cross-attention + MLP) maps output queries into the PDE solution. Our implementation is based on (Li et al., 2023a), but adapted for the 3D aerodynamics. [^9] https://github.com/BaratiLab/OFormer For our implementation, we use Galerkin attention. We use the standard ViT-tiny configuration with $12$ Transformer blocks, a channel size of $192$, $3$ self-attention heads, and an up-projection ratio of $4$.
Transolver
We reimplement Transolver according to the official implementation of (Wu et al., 2024) [^10] https://github.com/thuml/Transolver/blob/main/Car-Design-ShapeNetCar/models/Transolver.py and using a standard ViT-tiny configuration with $12$ Transolver blocks, a channel size of $192$, $3$ self-attention heads, and an up-projection ratio of $4$ for the feed-forward layers.. We use 512 slices in the Transolver attention, which we chose after varying the number of slices, where we did not obtain significant improvements for larger numbers of slices.
Transolver++
As Transolver++ (Luo et al., 2025) has no publicly available implementation, we attempted to reimplement it to the best of our abilities and tried to infer implementation details that are not outlined in the paper. However, contrary to the results of the original paper, our reimplementation performed consistently worse than the vanilla Transolver (Wu et al., 2024). Therefore, we do not consider it in the experiments section. We refer to Appendix C.2 for the full reproducibility ablation.
UPT
B.4.1 Data pre-processing
Input coordinates
Output targets
Input size
B.5 Training details
B.5.1 General configurations
For each baseline model, we used the AdamW (Loshchilov & Hutter, 2019) optimizer, with the exception of AB-UPT/UPT/Transformer, where we use LION (Chen et al., 2023). We tuned the learning rate of models trained with AdamW by sweeping over $\{1e^{-3},5e^{-3},2e^{-3},5e^{-3},1e^{-4},5e^{-5},2e^{-5}\}$ to achieve optimal performance per baseline. Whenever LION is used, we use $5e^{-5}$ as this value has performed consistently well. Training was conducted using either float16 or bfloat16 precision, depending on which yielded the best results. An exception was made for GINO, as it exhibited unstable behavior with float16/bfloat16 mixed precision. All models were trained for $500$ epochs with a batch size of $1$. We applied a weight decay of $0.05$ and a gradient clipping value of $1$. A cosine learning rate schedule was implemented, including a $5$% warm-up phase and a final learning rate of $1e^{-6}$
B.5.2 Model specific hyper-parameters
Per model, we tuned some model-specific details:
PointNet
LNO
Graph-UNet
GINO
UPT
Model LR Precision Model specific hyper-params Model Params
PointNet {$1e^{-3}$,$5e^{-4}$,$5e^{-4}$} float16 64 3.6M
GRAPH U-NET $1e^{-3}$float16, bfloat16, float16} - 14.1M
GINO {$1e^{-4}$,$2e^{-4}$,$5e^{-4}$} float32 - 15.6M
LNO {$5e^{-4}$,$2e^{-4}$,$1e^{-3}$} {bfloat16, float16, bfloat16} {2056, 256, 1024} 6.3M
UPT $5e^{-5}$ float16 {3586, 16384, 16384} 11.0M
OFormer {$2e^{-4}$,$2e^{-4}$,$1e^{-3}$} bfloat16 - 6.1M
Transolver $1e^{-3}$ bfloat16 - 5.5M
Transformer $5e^{-5}$ float16 - 5.5M
AB-UPT $5e^{-5}$ float16 - {6.7M, 6.7M, 8.8M}
B.5.3 Evaluation variance reduction
  1. (i)
  2. (ii)
  3. (iii)
  4. (iv)
Motivation
Note that this protocol was designed with limitations of baseline models in mind. AB-UPT can decode meshes of any size on a single GPU, train with batch size $>1$ on a single GPU (i.e., memory constraints are not the reason why we train with batch size 1), and decode more inputs during inference than in training.
Appendix C Baseline fairness and reproducibility
C.1 Transolver and baseline reproducibility

### v4

@@ 5 Limitations and future work
AB-UPT employs expressive self-attention among anchor tokens, which exhibits quadratic complexity. We show that on high-fidelity CFD simulations, a small set of anchor tokens is sufficient to obtain good results on currently available public datasets. However, using more anchor tokens can improve accuracy further, leading to (quadratically) increasing computational demands. Contrary, the computational demands of models with linear complexity would only increase linearly.
@@
Further improvements of AB-UPT and neural surrogates
While AB-UPT shows strong results, there are many interesting directions for further improvement of AB-UPT and neural surrogate models in general. Instead of randomly sampling anchor tokens, the sampling could be informed by domain expertise (e.g., dense sampling at boundary regions), uncertainty estimates, or even employ some kind of learned sampling procedure (e.g., by learning a probability distribution over the domain or using learnable 3D vectors as anchor positions). Additionally, there are many avenues to improve the performance of neural surrogates in general, including physics-preserving data augmentation, jointly training on high- and low-fidelity simulations or other heterogeneous data sources, transfer learning to low data regimes, low-fidelity to high-fidelity transfer learning, and uncertainty quantification to warn users of potentially high error regions or out-of-distribution settings. We believe many of these techniques could improve AB-UPT and neural surrogates in general; however, these are beyond the scope of this work.
@@
Fidelity vs variability trade-off
In this paper, we mainly consider high-fidelity HRLES CFD simulations, which can be extremely expensive to obtain (see Table 2). Consequently, current publicly available datasets tend to either produce high-fidelity simulations with relatively few input variability or lower-fidelity simulations with more input variability. For example, DrivAerML (Ashton et al., 2024b) uses high-fidelity HRLES simulations with a single geometry class and static simulation parameters (e.g., Reynolds number), whereas DrivAerNet++ (Elrefaie et al., 2024b) uses lower-fidelity RANS simulations with multiple geometry classes, sacrificing simulation fidelity in favor of geometric variability (simulation parameters are also static for DrivAerNet++). Evaluating AB-UPT on datasets with more diverse geometries and flow conditions is an interesting future direction.
Physics constraints via neural field formulation
In Section 4.7, we show how formulating a neural surrogate model as a conditional neural field (e.g., via anchor attention) can be leveraged to adhere to physics constraints such as predicting divergence-free vorticity by estimating differential operators via finite-differences (or automatic differentiation). We show this capability on a comparatively simple use-case of predicting divergence-free vorticity. Future work could leverage the neural field formulation to impose additional physics constraints, such as mass-conservation of the velocity field. However, this is a non-trivial extension as the curl operator alone is not sufficient to model a mass-conserving velocity field due to potential additive components that are simultaneously divergence-free and vorticity-free (i.e, harmonic).
@@ 6 Conclusion
Appendix A Model design ablations
A.1 Model design on DrivAerML
@@
A.2 Geometry branch design
@@
A.3 Training without query tokens
Training with loss on Loss when evaluated on
(i) 0.13574 0.13624 $0.37\%$ 11.5s
(ii) 0.13634 0.13632 $\mathbf{-0.01\%}$ 19.4s
(iii) 0.13462 0.13468 $\mathbf{+0.04\%}$ 19.4s
A.4 Variance of anchor selection
We evaluate the variance in predictive performance originating from the random sampling process of the anchor tokens. To evaluate this, we change the evaluation protocol of Section 4.4 (as described in Section C.5.3) slightly. Instead of propagating only a chunk of 16K anchor tokens (without any query tokens), then concatenating all chunks of a sample and calculating the relative L2 error on the concatenated chunks, we instead measure performance on query tokens, where we use the concatenation of all chunks as query positions. This results in multiple predictions for the exact same locations, where the only difference is the selection of the anchor points, where mean and standard deviation are calculated per sample and then averaged over the whole test split of the dataset. Table 11 shows that AB-UPT is fairly robust against the random anchor selection process.
Table 11: Relative L2 errors (in %) of surface pressure $\bm{p}_{s}$, volume velocity $\bm{u}$, volume vorticity $\bm{\omega}$, wall shear stress $\bm{\tau}$, and volume pressure $\bm{p}_{v}$ on the DrivAerML dataset. Lower percentage values indicate better performance in terms of L2 error. We evaluate the mean and standard deviation of a single seed of an AB-UPT model from Section 4.4 by fixing query locations and varying the anchor tokens.
$\bm{p}_{s}$ $\bm{u}$ $\bm{\omega}$ $\bm{\tau}$ $\bm{p}_{v}$
Variable anchors (from Section 4.4) 3.82 5.93 35.1 7.29 6.08
[…]
A.5 Anchor attention versus k-NN interpolation
Anchor attention updates the representation of a query token based on the anchor tokens, without influencing them. Additionally, we sample random points in space, so each query token most likely has a relatively close-by anchor token. To validate the effectiveness of anchor attention, we compare our approach against a baseline of a k-NN interpolation (Qi et al., 2017b), where we interpolate the anchor tokens after the full forward pass to query locations via a weighted sum of the 3 nearest neighbors of each query location, where the weights are based on the distance to each neighbor.
Table˜12 shows that anchor attention, although only trained on a reduced resolution (e. g., trained with 16K anchor points), can produce high-fidelity predictions also for query points (e. g.., up to 140 million points). It is significantly better than a simple k-NN interpolation as the model learned non-linear interactions between anchor points during training, which it can transfer “zero-shot”-like to query points, even though the query points do not influence the anchor points. This underlines the effectiveness of our approach and confirms that our model learns more sophisticated prediction schemes than linear interpolation.
@@ A.5 Anchor attention versus k-NN interpolation
$\bm{p}_{s}$ $\bm{u}$ $\bm{\omega}$
k-NN interpolation 0.00703 0.06641 0.5748
Anchor attention 0.00276 0.00657 0.0397
A.6 Mixed-precision training
We train all AB-UPT in fp16 mixed precision as this makes training and inference much faster. Previous works found instabilities when training with fp16 (Alkin et al., 2024a), which we also experienced in early development cycles. By keeping positional embeddings in fp32 precision and adding Rotary Positional Embeddings (RoPE) (Su et al., 2024) (also in fp32 precision), we stabilized training while preserving predictive accuracy and the large speedups of fp16 mixed precision training. To highlight this, we train AB-UPT in full fp32 precision on DrivAerML. Table 13 shows no significant predictive accuracy gain of fp32 training at vastly increased training times.
Due to the vast speedup of fp16 mixed precision training, we also train baselines in mixed precision whenever possible, where we also keep positional embeddings in fp32. For more detail,s see Appendix C.5.
Table 13: Relative L2 errors (in %) of surface pressure $\bm{p}_{s}$, volume velocity $\bm{u}$, volume vorticity $\bm{\omega}$, wall shear stress $\bm{\tau}$, and volume pressure $\bm{p}_{v}$ on DrivAerML for different training precisions. Training time is measured with 500 epochs on a single NVIDIA H100 GPU.
$\bm{p}_{s}$ $\bm{u}$ $\bm{\omega}$ $\bm{\tau}$ $\bm{p}_{v}$ Training time
AB-UPT (full fp32) 3.78 5.94 36.1 7.25 6.07 33.6h
AB-UPT (mixed-fp16) 3.82 5.93 35.1 7.29 6.08 6.9h
Appendix B Extended benchmark results
B.1 DrivAerNet++ evaluation
Figure 10: Relative improvement (in $\%$) of the test error (MSE) when increasing the model capacity via the number of surface and volume points. Test MSE improvement denotes the test MSE at the respective number of surface/volume points divided by the obtained test MSE at 2K points. Results are shown for Transolver, Transformer, and AB-UPT on both DrivAerNet++ and DrivAerML. Since DrivAerNet++ does not have vorticity, we exclude it from DrivAerML as well for better comparison. Increasing the model’s computational capacity leads to significant performance gains on DrivAerML, but has much smaller performance gains on DrivAerNet++.
Table 14: Relative L2 errors (in %) of surface pressure $\bm{p}_{s}$, wall shear stress $\bm{\tau}$, volume pressure $\bm{p}_{v}$, and volume velocity $\bm{u}$ on DrivAerNet++.
$\bm{p}_{s}$ $\bm{\tau}$ $\bm{p}_{v}$ $\bm{u}$
PointNet 21.62 31.76 20.07 16.75
GRAPH U-NET 18.68 27.42 16.13 13.62
GINO 15.45 23.74 16.24 13.80
LNO 21.87 31.93 18.00 15.68
UPT 14.31 22.16 12.55 11.06
OFormer 19.13 29.06 17.69 15.20
Transolver 13.97 21.73 12.44 11.06
Transformer 13.86 21.59 12.42 11.00
AB-UPT 13.79 21.49 12.34 10.94
We also evaluated AB-UPT on the DrivAerNet++ dataset (Elrefaie et al., 2024a; b), an open-source collection of over 8,000 automotive aerodynamic simulations with extensive geometric variations. While this dataset contains significantly more geometric variations than DrivAerML (Ashton et al., 2024b), a key limitation is its reliance on a lower-fidelity turbulence model, as the simulations were carried out using RANS. We compare AB-UPT against all baselines and use the same evaluation setup as in the main paper (see Section˜C.5.3). All models were trained to predict all fields at the same time as in our experiments in Section˜4.4, except that volume vorticity is missing from this dataset. For AB-UPT, super node pooling used a radius of $1.0$, the number of supernodes and the number of anchor points in the surface and volume branch was set to 16384 each, and, similarly to the main experiments (cf. Section˜C.3), the model uses 12 Transformer blocks, where the last $4$ self-attention blocks do not share parameters. Furthermore, we used a learning rate of $1\times 10^{-4}$ for AB-UPT and $5\times 10^{-5}$ for the baselines, and we trained all models for 35 epochs. Surface pressure $\bm{p}_{s}$, surface friction $\bm{\tau}$, volume pressure $\bm{p}_{v}$, and volume velocity $\bm{u}$ are reported in Table˜14. While AB-UPT outperforms all baselines, we notice that Transolver and Transformer follow much closer in performance compared to experiments on ShapeNet-Car, AhmedML, and DrivAerML in Table˜4.
Furthermore, it is evident that all surrogate models achieve worse performance in terms of relative L2 error on DrivAerNet++ than they do on DrivAerML or ShapeNet-Car. We attempted to train better models by, e.g., increasing training duration, increasing model size, or increasing the number of (anchor) points per model. In these experiments, we found that the training loss steadily declines where AB-UPT obtains significantly lower train losses than Transolver and a vanilla Transformer, while test performance (measured by relative L2 errors) does not improve with any of the three considered models. We also observed that increasing model capacity leads to significantly less performance gains on DrivAerNet++ compared to DrivAerML (see Figure˜10). This suggests that model capacity is not the limiting factor, but there could be significant irreducible noise in the dataset resulting from the lower-fidelity simulation method. Determining the primary cause of this performance discrepancy was not investigated further, but it would be a valuable direction for future research.
B.2 Benchmarking AB-UPT against TripNet on DrivAerNet++
We compare AB-UPT against the recently proposed TripNet (Chen et al., 2025) on the DrivAerNet++ dataset (Elrefaie et al., 2024a; b). As there is no publicly available implementation of TripNet and the input representation and pre-processing pipeline is completely different from our setup, we did not attempt to implement this baseline and instead compare directly to the results reported in Chen et al. (2025). We trained AB-UPT with the train/validation/test split provided by Elrefaie et al. (2024b) and report median evaluation metrics over 5 training runs. For evaluating the vector-valued fields, we follow the experimental setup of Chen et al. (2025) and report the relative L2 error of the wall shear stress magnitude $|\bm{\tau}|$ and velocity magnitude $|\bm{u}|$. The results are summarized in in Table˜15, and show that AB-UPT significantly outperforms TripNet on all metrics.
Table 15: Relative L2 errors (in %) of surface pressure $\bm{p}_{s}$, surface friction wall shear stress magnitude $|\bm{\tau}|$, and volume velocity magnitude $|\bm{u}|$, and the velocity components in the respective directions. Lower percentage values indicate better performance in terms of L2 error. We provide the results for DrivAerNet++. AB-UPT outperforms TripNet, often by quite a margin. The results for TripNet are taken from (Chen et al., 2025)*.
Surface Volume
$\bm{p}_{s}$ $|\bm{\tau}|$ $|\bm{u}|$ $\bm{u}_{x}$ $\bm{u}_{y}$ $\bm{u}_{z}$
TripNet* 20.05 22.07 10.39 10.71 35.34 36.39
AB-UPT 13.79 18.26 8.48 8.63 32.18 30.98
B.3 Benchmarking AB-UPT against DoMINO
We trained AB-UPT with the same train/test split as provided in (Ranade et al., 2025), and report median evaluation metrics over 5 training runs. All other hyperparameters for AB-UPT remained consistent with those reported in Appendix C. The numbers for DoMINO are taken directly from (Ranade et al., 2025) and we compare to the same metrics.
Table 16: Relative L2 errors (in %) of surface pressure $\bm{p}_{s}$, wall shear stress $\bm{\tau}$ per $x,y,z$, dimension, volume velocity $\bm{u}$ per $x,y,z$ dimension, and volume pressure $\bm{p}_{v}$. Lower percentage values indicate better performance in terms of L2 error. We provide the results for DrivAerML. AB-UPT outperforms DoMINO, often by quite a margin. The results for DoMINO are taken from (Ranade et al., 2025)*.
Surface Volume
$\bm{p}_{s}$ $\bm{\tau}_{x}$ $\bm{\tau}_{y}$ $\bm{\tau}_{z}$ $\bm{p}_{v}$ $\bm{u}_{x}$ $\bm{u}_{y}$ $\bm{u}_{z}$
DoMINO* 15.05 21.24 30.2 33.59 21.93 23.97 50.25 45.67
AB-UPT 3.76 5.35 3.65 3.63 6.29 4.43 3.04 2.61
B.4 Benchmarking AB-UPT against Erwin
Table 17: Benchmarking our adaptation of Erwin to jointly model surface and volume data on ShapeNet-Car. Relative L2 errors (in %) of surface pressure $\bm{p}_{s}$, and volume velocity $\bm{u}$. Erwin (reimpl.) with ball size 8192 would be roughly equivalent to the Transformer row.
Ball size $\bm{p}_{s}$ $\bm{u}$
Linear complexity Transformers
LNO - 9.05 2.29
OFormer - 7.05 1.60
Transolver - 6.46 1.62
Erwin (reimpl.) 512 5.46 1.37
Erwin (reimpl.) 1024 5.38 1.37
Erwin (reimpl.) 2048 5.35 1.27
Erwin (reimpl.) 4096 5.31 1.25
Quadratic complexity Transformers
Transformer 8192 4.86 1.17
AB-UPT 4096 4.81 1.16
Next, we compare our reimplementation of Erwin on DrivAerML in the default setting considered in the main paper (16K/16K surface/volume points) and show results in Figure 11. We observe that ball attention imposes an overhead, most likely from memory restructuring operations to structure tokens for efficient attention computations. This overhead becomes negligible with more inputs, and ball attention becomes faster than (branched) self-attention. As there is a significant accuracy gap between AB-UPT trained on 16K/16K points and Erwin (reimpl.) trained on 16K/16K points, we hypothesize that more sophisticated training protocols are necessary to train ball tree attention models to sufficient quality (e.g., hierarchical architectures, different ball size to input token ratios, etc.). As the design space thereof is large and fair comparisons to AB-UPT become more and more difficult, we leave exploration of this direction to future work.
Finally, we compare AB-UPT against the original Erwin architecture (including three-layer message passing and distance-based attention biases) in the setting used in Zhdanov et al. (2025). Table 18 confirms that our reimplementation obtains sensible results and that AB-UPT also performs best with this protocol.
$\bm{p}_{s}$ MSE
Erwin-S 15.85
Erwin-M 15.43
Erwin (reimpl.) 15.17
AB-UPT 13.25
B.5 Extended comparison against Transolver
In Transolver (Wu et al., 2024), the quadratic attention is limited to a subset of tokens by slicing the input domain into a reduced domain (i.e., physics tokens). Throughout the experiments in the main paper, we use 512 slices. In this section, we investigate the impact of this hyperparameter in terms of performance (test loss), memory consumption, and runtime. Figure 12 shows that increasing the number of slices slightly improves test loss at the cost of additional memory and runtime requirements. As the slicing mechanism is implemented via slow and memory-heavy torch.einsum operations (following the original implementation [^5] https://github.com/thuml/Transolver/blob/main/Car-Design-ShapeNetCar/models/Transolver.py), it requires much more memory than AB-UPT, which leverages the constant memory complexity of FlashAttention (Dao et al., 2022) (which results in constant memory complexity for the attention layer, but the overall memory complexity is still linear). For larger problem sizes (e.g., DrivAerML), the memory bottleneck becomes significant and requires memory-saving techniques such as activation checkpointing to fit onto a single GPU, which in turn increases runtime. Note that we improved speed and memory consumption of the original Transolver implementation by using FlashAttention for the self-attention between slice tokens and employing mixed-precision training. Further optimizations (e.g., specialized hardware implementations of the transolver attention) are certainly possible but are beyond the scope of this work. Note that we also do not optimize AB-UPT beyond the use of FlashAttention666torch.nn.functional.scaled_dot_product_attention using PyTorch 2.4.1 with CUDA 12.4 with mixed-precision training. Also AB-UPT could be optimized further, e.g., by employing FlashAttention-3 (Shah et al., 2024).
B.6 Transolver and baseline reproducibility